OpenStudy (anonymous):

HELP! prove tan5utan3u=(tan^2*4u-tan^2*u)/(1-tan^2*4u*tan^2u

hartnn (hartnn):

you know tan(A+B) and tan(A-B) formula ?? write 5u = 4u+u 3u=4u-u and use the formula.

hartnn (hartnn):

\[\tan(A+B) = \frac{\tan A+\tan B}{1-\tan A.\tan B}\] put, A=4u, B=u here. \[\tan(A-B) = \frac{\tan A+\tan B}{1+\tan A.\tan B}\] put, A =4u, B =u here also, then just multiply both the equations, and you'll get your result.